OpenStudy (anonymous):

When writing linear equations, how do you determine which form of a line to use? Part 2: Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. Point pairs (5, 1), (–3, 4) (0, –2), (3, 2) (–2, –1), (1, 2)

OpenStudy (solomonzelman):

idk, this question has to do with the content of your lesson, because I (almost) always prefer slope intercept formula.

OpenStudy (anonymous):

To me, although equation of a line has 2-3 forms, but they are derived from only 1 standard form: y = mx +b

OpenStudy (anonymous):

for part2) you have to: 1/ find the slope of the line by using the formula m = \(\dfrac{y_2-y_1}{x_2-x_1}\) 2/ pick a point (either of them) and plug the coordinate into the formula y = mx +b